Use the at least once rule to find the probabilities of the following events?

1.) Getting at least one tail when tossing five fair coins?

2.) Getting at least one 50 year flood in the next ten years?
Answer #1

solution - þ ( tails) = 1/2 り P(at least one tails in tossing stair coins) 1 - Pimone Hails) 1 (1-P) = 1-11-16) 0.96875 บ plt